# Interplay of Zeeman field, Rashba spin-orbit interaction, and superconductivity: Transition temperature and quasiparticle excitations

### v1

@@ Appendix F Bogoliubov transformation: $p$-wave pairing states in a single Zeeman field
$\begin{split}&U_{\mathbf{k}}=U_{1}U_{2\mathbf{k}}\\ &=\begin{pmatrix}\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}&0&0\\ \sin{\frac{\theta}{2}}e^{i\phi}&-\cos{\frac{\theta}{2}}e^{i\phi}&0&0\\ 0&0&\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}\\ 0&0&\sin{\frac{\theta}{2}}e^{-i\phi}&-\cos{\frac{\theta}{2}}e^{-i\phi}\end{pmatrix}\cdot\frac{1}{2\sqrt{E_{\mathbf{k}\parallel}}}\\ &\begin{pmatrix}s_{\mathbf{k}3}e^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-s_{\mathbf{k}3}e^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&s_{\mathbf{k}3}e^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-s_{\mathbf{k}3}e^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&-e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-e^{i\gamma_{\mathbf{k}}}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ -s_{\mathbf{k}3}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&s_{\mathbf{k}3}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&s_{\mathbf{k}3}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-s_{\mathbf{k}3}e^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ \sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\end{pmatrix},\end{split}$ (85)
@@
$\begin{split}d_{12}(\mathbf{k})&=|\mathbf{d}(\mathbf{k})\times\hat{\mathbf{H}}|,\\ d_{12}(\mathbf{k})e^{i\Omega_{\mathbf{k}}}&=d_{1}(\mathbf{k})+id_{2}(\mathbf{k}),\\ s_{\mathbf{k}3}&=\mbox{sgn}[d_{3}(\mathbf{k})].\end{split}$ (86)
@@
$\begin{split}u_{\mathbf{k}}&=\frac{e^{i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\cos\frac{\theta}{2}+e^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\sin\frac{\theta}{2}&e^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\sin\frac{\theta}{2}-s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\cos\frac{\theta}{2}\\ e^{i\phi}\left(s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\sin\frac{\theta}{2}-e^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\cos\frac{\theta}{2}\right)&-e^{i\phi}\left(e^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\cos\frac{\theta}{2}+s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\sin\frac{\theta}{2}\right)\end{pmatrix},\\ v_{\mathbf{k}}&=\frac{1}{2}\begin{pmatrix}a_{\mathbf{k}1+}b_{\mathbf{k}-}\sin\frac{\theta}{2}-e^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\cos\frac{\theta}{2}&e^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\cos\frac{\theta}{2}+a_{\mathbf{k}2-}b_{\mathbf{k}+}\sin\frac{\theta}{2}\\ -e^{i\phi}\left(a_{\mathbf{k}1+}b_{\mathbf{k}-}\cos\frac{\theta}{2}+e^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\sin\frac{\theta}{2}\right)&e^{i\phi}\left(e^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\sin\frac{\theta}{2}-a_{\mathbf{k}2-}b_{\mathbf{k}+}\cos\frac{\theta}{2}\right)\end{pmatrix},\\ v_{-\mathbf{k}}^{*}&=\frac{1}{2}\begin{pmatrix}a_{\mathbf{k}1+}b_{\mathbf{k}-}\sin\frac{\theta}{2}-e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\cos\frac{\theta}{2}&e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\cos\frac{\theta}{2}+a_{\mathbf{k}2-}b_{\mathbf{k}+}\sin\frac{\theta}{2}\\ -e^{-i\phi}\left(a_{\mathbf{k}1+}b_{\mathbf{k}-}\cos\frac{\theta}{2}+e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\sin\frac{\theta}{2}\right)&e^{-i\phi}\left(e^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\sin\frac{\theta}{2}-a_{\mathbf{k}2-}b_{\mathbf{k}+}\cos\frac{\theta}{2}\right)\end{pmatrix},\\ u_{-\mathbf{k}}^{*}&=-\frac{e^{-i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\cos\frac{\theta}{2}+e^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\sin\frac{\theta}{2}&e^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\sin\frac{\theta}{2}-s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\cos\frac{\theta}{2}\\ e^{-i\phi}\left(s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\sin\frac{\theta}{2}-e^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\cos\frac{\theta}{2}\right)&-e^{-i\phi}\left(e^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\cos\frac{\theta}{2}+s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\sin\frac{\theta}{2}\right)\end{pmatrix}.\end{split}$ (88)
@@
Appendix G Bogoliubov transformation: $p$-wave pairing states with only Rashba SOC

### v2

@@ Appendix G Bogoliubov transformation: $p$-wave pairing states in a single Zeeman field
$\begin{split}&U_{\mathbf{k}}=U_{1}U_{2\mathbf{k}}\\ &=\begin{pmatrix}\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}&0&0\\ \sin{\frac{\theta}{2}}\mathrm{e}^{i\phi}&-\cos{\frac{\theta}{2}}\mathrm{e}^{i\phi}&0&0\\ 0&0&\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}\\ 0&0&\sin{\frac{\theta}{2}}\mathrm{e}^{-i\phi}&-\cos{\frac{\theta}{2}}\mathrm{e}^{-i\phi}\end{pmatrix}\cdot\frac{1}{2\sqrt{E_{\mathbf{k}\parallel}}}\\ &\begin{pmatrix}s_{\mathbf{k}3}\mathrm{e}^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-s_{\mathbf{k}3}\mathrm{e}^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&s_{\mathbf{k}3}\mathrm{e}^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-s_{\mathbf{k}3}\mathrm{e}^{i\gamma_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ \mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&-\mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-\mathrm{e}^{i\gamma_{\mathbf{k}}}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ -s_{\mathbf{k}3}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&s_{\mathbf{k}3}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&s_{\mathbf{k}3}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&-s_{\mathbf{k}3}\mathrm{e}^{i\Omega_{\mathbf{k}}}\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\\ \sqrt{\frac{(E_{\mathbf{k}+}+E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}-E_{\mathbf{k}\parallel}+\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}&\sqrt{\frac{(E_{\mathbf{k}+}-E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}-\xi_{\mathbf{k}})}{E_{\mathbf{k}+}}}&\sqrt{\frac{(E_{\mathbf{k}-}+E_{\mathbf{k}\parallel}-\mu_{B}|\mathbf{H}|)(E_{\mathbf{k}\parallel}+\xi_{\mathbf{k}})}{E_{\mathbf{k}-}}}\end{pmatrix},\end{split}$ (85)
@@
$\begin{split}d_{12}(\mathbf{k})&=|\mathbf{d}(\mathbf{k})\times\hat{\mathbf{H}}|,\\ d_{12}(\mathbf{k})\mathrm{e}^{i\Omega_{\mathbf{k}}}&=d_{1}(\mathbf{k})+id_{2}(\mathbf{k}),\\ s_{\mathbf{k}3}&=\mbox{sgn}[d_{3}(\mathbf{k})].\end{split}$ (86)
@@
$\begin{split}u_{\mathbf{k}}&=\frac{\mathrm{e}^{i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\cos\frac{\theta}{2}+\mathrm{e}^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\sin\frac{\theta}{2}&\mathrm{e}^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\sin\frac{\theta}{2}-s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\cos\frac{\theta}{2}\\ \mathrm{e}^{i\phi}\left(s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\sin\frac{\theta}{2}-\mathrm{e}^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\cos\frac{\theta}{2}\right)&-\mathrm{e}^{i\phi}\left(\mathrm{e}^{i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\cos\frac{\theta}{2}+s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\sin\frac{\theta}{2}\right)\end{pmatrix},\\ v_{\mathbf{k}}&=\frac{1}{2}\begin{pmatrix}a_{\mathbf{k}1+}b_{\mathbf{k}-}\sin\frac{\theta}{2}-\mathrm{e}^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\cos\frac{\theta}{2}&\mathrm{e}^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\cos\frac{\theta}{2}+a_{\mathbf{k}2-}b_{\mathbf{k}+}\sin\frac{\theta}{2}\\ -\mathrm{e}^{i\phi}\left(a_{\mathbf{k}1+}b_{\mathbf{k}-}\cos\frac{\theta}{2}+\mathrm{e}^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\sin\frac{\theta}{2}\right)&\mathrm{e}^{i\phi}\left(\mathrm{e}^{-i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\sin\frac{\theta}{2}-a_{\mathbf{k}2-}b_{\mathbf{k}+}\cos\frac{\theta}{2}\right)\end{pmatrix},\\ v_{-\mathbf{k}}^{*}&=\frac{1}{2}\begin{pmatrix}a_{\mathbf{k}1+}b_{\mathbf{k}-}\sin\frac{\theta}{2}-\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\cos\frac{\theta}{2}&\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\cos\frac{\theta}{2}+a_{\mathbf{k}2-}b_{\mathbf{k}+}\sin\frac{\theta}{2}\\ -\mathrm{e}^{-i\phi}\left(a_{\mathbf{k}1+}b_{\mathbf{k}-}\cos\frac{\theta}{2}+\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}2+}b_{\mathbf{k}+}\sin\frac{\theta}{2}\right)&\mathrm{e}^{-i\phi}\left(\mathrm{e}^{i\Omega_{\mathbf{k}}}s_{\mathbf{k}3}a_{\mathbf{k}1-}b_{\mathbf{k}-}\sin\frac{\theta}{2}-a_{\mathbf{k}2-}b_{\mathbf{k}+}\cos\frac{\theta}{2}\right)\end{pmatrix},\\ u_{-\mathbf{k}}^{*}&=-\frac{\mathrm{e}^{-i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\cos\frac{\theta}{2}+\mathrm{e}^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\sin\frac{\theta}{2}&\mathrm{e}^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\sin\frac{\theta}{2}-s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\cos\frac{\theta}{2}\\ \mathrm{e}^{-i\phi}\left(s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}\sin\frac{\theta}{2}-\mathrm{e}^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}2+}b_{\mathbf{k}-}\cos\frac{\theta}{2}\right)&-\mathrm{e}^{-i\phi}\left(\mathrm{e}^{-i\Omega_{\mathbf{k}}}a_{\mathbf{k}1-}b_{\mathbf{k}+}\cos\frac{\theta}{2}+s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\sin\frac{\theta}{2}\right)\end{pmatrix}.\end{split}$ (88)
@@
Appendix H Bogoliubov transformation: $p$-wave pairing states with only Rashba SOC